Biochemistry: Please make the growth curve on excel or throughly explain the solution so I can make the graph and how you did it.
minutes es ominues → 0.500-→ 0.uo5 SroeT 15 minutr DISCUSSION o.3叫ー→0.430- I. Make a growth curve of your E. coli culture that was induced with IPTG. Calculate the doubling time of your culture. How does your doubling time compare to the optimal doubling time of E. coli and how do you think this influenced protein yield?
Answer #1

1. take od on x axis and time on y axis and make a curve by joining the lines thus u get a growth curve of ecoli

doubling time g=1/k

k=n/t [ mean growth rate constant (k)]

n is number of generations

t is total time taken

k=3/35g=1/k  

g=35/3

=11.6

mean doubling time of ecoli is 20min which is almost double to our value[11.6]

this will lead todoubles the protein yield .